# On Leakage of Code Generation Evaluation Datasets

### v2

@@ Abstract
In this paper we consider contamination by code generation test sets, in particular in their use in modern large language models. We discuss three possible sources of such contamination and show findings supporting each of them: (i) direct data leakage, (ii) indirect data leakage through the use of synthetic data and (iii) overfitting to evaluation sets during model selection.
Key to our findings is a new dataset of 161 prompts with their associated python solutions, dataset which is released at https://huggingface.co/datasets/CohereForAI/lbpp.
@@ 1 Introduction
Code generation has emerged as an important skill for large language models to master. Measuring recent progress in code generation has relied on few, critical benchmarks to judge performance between model families and checkpoints. While many recent sophisticated evaluation datasets have been proposed (Jain et al., 2024; Jimenez et al., 2024), the community largely relies on HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) to judge a new model’s code capability. In fact, all major announcements in 2023-2024 claiming advanced code capabilities—from academic and industry labs—use at least one of these two datasets. Practically, reporting HumanEval and MBPP is mandatory for a model to report competitive code generation.
@@
However, the importance of these benchmarks has led to a conflict between popularity and utility. On one side, obtaining competitive numbers comes with significant scientific and economic reward—made increasingly easy with the proliferation of public replicas of these datasets. However, this prevalence has led to data leakage beyond the original evaluation scope, i.e., data contamination, and once this evaluation data contaminates model training, the validity of the metrics as a measure of generalization capability becomes unreliable. If a model has been trained on the same data we use for out-of-distribution generalization (or is selected based on its performance on that data), we break an implicit tenet of how model capability can be measured. We argue that understanding the effect of contamination is critical to accurately interpreting scores on these benchmarks.
@@
In this paper, we review the evidence that these two benchmarks have contaminated most large LLMs, which we define as any procedure that leaked those datasets during model training. The most obvious method of contamination is presence inside training data, and we provide evidence that it is highly probably that this occurs at a scale too large to be avoidable. A second possibility is that contamination happens indirectly through the use of synthetic data—a widespread paradigm used in particular to increase code capabilities by generating additional code training tokens. Finally, we argue that final model selection might have been overly influenced by their performance on these datasets, overfitting to performance on these metrics over general-purpose code-oriented skills.
To measure this contamination, we propose Less Basic Python Problems (LBPP), a code generation benchmark similar to HumanEval and MBPP in style and scale, but more difficult. LBPP is similarly portable, but is produced in a manner to reduce any likelihood of leakage into present code training data. We contribute LBPP to act as a genuinely held-out test set to measure current code generation capability, and potential overfitting to HumanEval and MBPP.
@@ 2 Related Work
HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) remain the most reported results on public leaderboards, but others similar datasets exist (Hendrycks et al., 2021; Li et al., 2022). They consist of short and mostly simple (not programming competition level) instructions with completions in Python. Translation into other programming languages exist for those datasets (Muennighoff et al., 2023; Cassano et al., 2022), as well as versions with additional tests (Liu et al., 2024).
@@
(Jain et al., 2024) proposed a continuously updated set of leetcode to improve dataset challenge by including harder and novel (unseen) prompts. (Jimenez et al., 2024) aims for challenging software engineering problems, that require understanding of full repositories. In a similar vein, RepoQA [^1] https://github.com/evalplus/repoqa and Bug In The Code Stack [^2] https://github.com/HammingHQ/bug-in-the-code-stack focus on understanding long contexts within code tasks. One proposed solution is to use hidden evaluation sets (Zhang et al., 2024), however, these do not allow inspection of failure cases and requires trusting the quality and correctness of an opaque ‘black-box’ evaluation setup. Recently, Riddell et al. (2024) analyzed data contamination in popular pretraining datasets: reporting that $12.2\%$ of HumanEval samples are present in The Pile (Gao et al., 2020), and $18.9\%$ in The Stack (Kocetkov et al., 2022). Different from our analysis however, they conclude: “we do not find the performance rankings of the models to change with decontaminated results”.
@@
Figure 1: Pass@1 rate of popular datasets and the 161 prompts in LBPP.
3 Possible sources of contamination
We provide three hypotheses—with evidence for each one—on why existing models might be over-optimized towards existing leaked benchmarks.
3.1 Direct data leakage
The most obvious reason is the simplest: many of the test datasets are of widespread use and the simplest answer might be that modern LLMs are just trained on this evaluation data. We note that intentional (i.e., to cheat) or unintentional contamination has the same net effect: training on evaluation data limits the confidence and utility of the benchmark results. For code tasks, it is very expensive to curate datasets of natural language to code instructions (one example generally costing several dozen US dollars). For any group aiming to minimize data cost to improve coding performance, the dollar value of creating new datasets can be very high. This leads to a common practice of web scraping code-oriented resources (e.g., GitHub or Stackoverflow) for data. However, these resources are also likely sources of contamination. Since the release of HumanEval and MBPP in 2021 these datasets have been branched, re-used and copied all across the Internet. The small data size and portability of such benchmarks encourages replication within code repositories. For example, searching for the prompts from HumanEval on GitHub returns a hit in all cases—the median hits is $99$ and the minimum $43$ (see Fig. 2). In many cases, these hits are exact duplicates and indications of a fork of the original dataset.
While decontamination of training sets is becoming more common, present decontamination filters designed for natural text adapts poorly to code. To operate efficiently at scale, most filters rely on generic deduplication algorithms e.g., such as $n$-gram matching or hashing functions Lee et al. (2022). Such surface-level matching does not adequately capture code similarity where a simple variable name change leaves program semantics unchanged, but changing a single keyword can have profound changes. [^3] E.g., compare the instruction “return true if the string is a float” with “return true if the string is a verb”. The same shortcomings of decontamination efforts apply to the creation of large-scale synthetic datasets: for example the model-generated dataset of Starcoder (Li et al., 2023) is decontaminated only by removing exact docstrings or solutions that match HumanEval or MBPP.
The recent exploration of Riddell et al. (2024) aims to quantify the proportion of this data leakage in existing datasets using plagiarism tools specifically designed for code. Even when static training datasets are cleaned, contamination may persist. Entities who serve models through an API may encounter these benchmark tasks when evaluated by third party users. When a sample of real model usage is annotated for future training data, samples from benchmark evaluation can leak into future training corpora. Furthermore, these samples may include subtle phrasing variations and format changes that further complicate heuristic deduplication. In this scenario, a model may easily memorize completions to purportedly novel prompts. As evidence of this phenomena, we prompted one popular commercial system with partial prompts from HumanEval that were designed to keep the instruction under-specified. Table Appendix in the Appendix shows the outcome and evidence that—despite the ambiguity of the prompt—the resulting completion matches exactly the gold solution from the test set of HumanEval.
3.2 Data leakage through synthetic data
The most capable of code language models rely heavily on the use of synthetic training data (Xu et al., 2023; Wei et al., 2023, 2024).
A typical pipeline generally consists of: curating prompts related to code generation, inferring completions with a previously trained LLM, and synthesizing unit tests for relevant prompts using LLMs. Completions that pass the respective unit tests are considered valid code solutions and can be used as future training examples. Alternatively, if a sufficiently powerful model is used, completions might be used as-is.
evol-instruct for example comprises $110k$ complex query prompts coupled with completions from numerous closed and open-source models. [^4] Per downloads, the most popular version is a ‘lightly decontaminated’ version on HuggingFace here. It is widely used by many code LLMs such as WizardCoder (Xu et al., 2023). Prior reports  (Yu et al., 2023, page 8)(Wei et al., 2023, page 4) discuss an apparent high similarity between some of those examples in evol-instruct and HumanEval. We extend this analysis by studying the similarity between embedded representations of the prompts [^5] Embedded using Cohere embed v3 (Team, 2024). of HumanEval and MBPP with nearest neighbors from evol-instruct and Starcoder-Instruct. Fig 3 highlights widespread similarity between evaluation data and synthetic training datasets ‘Starcoder V2 OSS Instruct’ and ‘evol instruct’. Even if unintentional, this contamination further damages the utility of these benchmarks as a held-out test of code generation capability.
Strikingly, we saw that training on this dataset can increase performance of one of our models by 14 absolute points on HumanEval (from 0.52 pass@1 to 0.66), while only by 1 point on MBPP (from 0.52 to 0.53). Inspecting closest neighbours for almost all examples in HumanEval (see examples in Table Appendix in the Appendix)—we identify a semantically equivalent version in evol-instruct.
Even when the original evaluation datasets were not used as inspiration for the creation of the synthetic datasets the simple nature of the evaluation datasets might make duplication unavoidable. There are only so many short natural language prompts describing typical interview-style programming questions that can be used. While synthetically generated prompts may not be explicitly based on provided examples from a given test set in the prompt context, the massive scale of these datasets ($238k$ instances for StarCoder-Instruct prior to deduplication) runs the risk of exhausting the possible number of variations on questions. Table 1 shows examples where generated prompts are extremely similar to questions in the MBPP test set. Note that despite that similarity training on this dataset did not improve substantially performance on MBPP.
3.3 Overfitting to test sets
The exaggerated importance of these benchmarks encourages an incentive structure where model selection prioritizes gain on a narrow suite of metrics. While it may be tempting to use such benchmarks as a deciding factor between similar checkpoints, there is weak evidence for these benchmarks correlating with ‘solving code generation’. While the meaning and measurement of this unscientific objective is subject to constant revision, selecting for optimal HumanEval performance may be akin to $p$-hacking in other fields. This practice can be justified by assuming that these benchmarks are the new dev sets, while the true test is the usage of users over time. The risk remains however that some models overfit to those test sets more than others, distorting perception on relative performance.
In order to measure this, we created Less Basic Python Problems (LBPP), a dataset of 161 code completion problems in the style of HumanEval. Human annotators were instructed to create totally fresh problems, which were not solvable by an internal model [^6] We will update references to internal models in future revisions. they had access to. Annotators had competitive programming experience and could use programming books as inspiration, but were instructed not to copy existing solutions on the Internet and not to use any LLMs. All annotators were paid above minimum wage in their respective countries, and all final prompt-completion pairs were manually reviewed by the authors. This adversarial collection resulted in more difficult problems, with most models solving less than $50\%$ of the dataset. Results on a selection of models are in Table 4 in the Appendix.
Using this dataset we can correlate the performances of existing models on them against performance of the well-known benchmarks. The two plots comparing them to HumanEval and MBPP can be seen in Fig. 1. There is a clear correlation on both data-sets, indicating that the public benchmarks are still a valuable target signal. However, when zooming in — in particular for models from the same family — the correlation often becomes negative, which might indicate that the selected checkpoint to release performs better on those public datasets while under-performing on new datapoints. Note in particular in Fig. 1(a) the crowded space between 0.75 and 0.8 of the $x$-axis (pass@1 of HumanEval). Despite very similar performance on that public evaluation set, the numbers of LBPP vary wildly. On the other side, in Fig. 1(b) a vertical line can be traced at around $0.38$ of LBPP that crosses various points – indicating maybe an overfit to MBPP.

### v3

@@ Abstract
In this paper, we consider contamination by code generation test sets, in particular in their use in modern large language models. We discuss three possible sources of such contamination and show findings supporting each of them: (i) direct data leakage, (ii) indirect data leakage through the use of synthetic data and (iii) overfitting to evaluation sets during model selection. To address this, we release LBPP: an uncontaminated new benchmark of 161 prompts with their associated Python solutions. LBPP is released at https://huggingface.co/datasets/CohereForAI/lbpp.
@@ 1 Introduction
Code generation has emerged as an important skill for large language models to master. Measuring recent progress in code generation has relied on few, critical benchmarks to judge performance between model families and checkpoints. While many recent sophisticated evaluation datasets have been proposed (Jain et al., 2024; Jimenez et al., 2024), the community largely relies on HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) to judge a new model’s code capability. In fact, all major announcements in 2023-2024 claiming advanced code capabilities—from either academic or industry labs—boast one or both of these datasets. Practically, reporting HumanEval and MBPP is mandatory for a model to report competitive code generation.
@@
However, the importance of these benchmarks has led to a conflict between popularity and utility. Obtaining competitive numbers comes with significant scientific and economic reward—made increasingly easy with the proliferation of public replicas and evaluation harnesses featuring these datasets. However, this prevalence has led to data leakage beyond the original evaluation scope. When this evaluation data contaminates model training, the validity of the metrics as a measure of generalization capability becomes unreliable. If a model is trained on data used for out-of-distribution generalization (or selected for its performance on that data), we break an implicit tenet of how model capability should be measured. We argue that understanding the effect of contamination is critical to accurately interpreting scores on these benchmarks.
@@
In this paper, we review the evidence that most contemporary LLMs are contaminated with data sourced from these two benchmarks. We define contamination here as any procedure leaking datasets during model training such that these datasets are no longer unseen at inference. The most obvious method of contamination is the presence inside training data. Section 3.1 reviews evidence that these benchmarks are widespread in training corpora in original and paraphrased forms. Unfortunately, it is not feasible to manually remove all the corresponding examples from the training corpora and the most common automatic decontamination methods have low recall. Section 3.2 proposes that contamination also occurs indirectly through the use of synthetic data—a widespread paradigm used to increase coding capabilities by generating additional code training tokens for pre-training or fine-tuning. Finally, Section 3.3 argues that checkpoint selection may be overly influenced by these datasets, overfitting to these benchmarks over general-purpose code-oriented generalization.
In this paper, we propose a more challenging Python code generation benchmark: Less Basic Python Problems. LBPP is similar in size to HumanEval and MBPP, but designed to be more complex using model in the loop filtering. LBPP is also designed to share no inspiration or sources with existing training and evaluation datasets, presenting a novel generalization challenge to contemporary LLMs. In Section 4, we observe that SOTA models on HumanEval perform up to $43\%$ worse on LBPP. We contribute LBPP as a genuinely held-out test to measure current code generation capability, and potential overfitting to HumanEval and MBPP.
@@ 2 Related Work
HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) remain the most reported results on public leaderboards, but similar datasets exist (Hendrycks et al., 2021; Li et al., 2022). They consist of short and mostly simple (not programming competition level) instructions with completions in Python. These datasets have also been translated into more programming languages (Muennighoff et al., 2023; Cassano et al., 2022), as well as versions with additional tests (Liu et al., 2024b).
@@
Jain et al. (2024) proposes a continuously updated set of interview questions to improve dataset challenge by including harder and novel (unseen) prompts. Jimenez et al. (2024) aims for challenging software engineering problems requiring understanding of entire repositories. In a similar vein, RepoQA (Liu et al., 2024a) and Bug In The Code Stack (Lee et al., 2024) focus on understanding long contexts within code tasks. Zhang et al. (2024) proposes using hidden evaluation sets, however, this approach does not allow inspection of failure cases and requires trusting the quality and correctness of an opaque ‘black-box’ evaluation setup.
Recently, Riddell et al. (2024) analyzed data contamination in popular pretraining datasets: reporting that $12.2\%$ of HumanEval samples are present in The Pile (Gao et al., 2020), and $18.9\%$ in The Stack (Kocetkov et al., 2022). While Riddell et al. (2024) reports “we do not find the performance rankings of the models to change with decontaminated results”, we identify the ranking between models to vary between contaminated and uncontaminated evaluation datasets (see Table 2).
3 Possible sources of contamination
We provide three hypotheses, with respective evidence, on why existing evaluation datasets are leaked and models may already be over-optimized towards existing leaked benchmarks.
3.1 Direct data leakage
The most obvious reason is the simplest: many of the test datasets are of widespread use and the simplest answer is that modern LLMs are trained on this evaluation data. We note that intentional (i.e., to cheat) or unintentional contamination has the same net effect: training on evaluation data limits the confidence and utility of the benchmark.
Curating high-quality datasets of natural language to code instructions can incur exorbitant costs when one example may cost upwards of dozens of US dollars. For any party considering the Pareto optimality of dataset size and coding performance gain, the required funding to create novel data can quickly explode. This leads to a common practice of web scraping code-oriented resources (e.g., GitHub or Stack Overflow) for data. However, these resources are also likely sources of contamination. The small data size and portability of such benchmarks encourages replication. We demonstrate this proliferation by keyword searching for HumanEval prompts on GitHub. Fig. 1 shows that we return a hit in all cases—the median hits is $99$ and the minimum $43$. These hits are often exact duplicates indicating a fork of the original dataset.
While decontamination of training sets is becoming more common, present decontamination filters designed for natural text adapt poorly to code. To operate efficiently at scale, most filters rely on generic deduplication algorithms e.g., such as $n$-gram matching or hashing functions Lee et al. (2022). Such surface-level matching does not adequately capture code similarity where a simple variable name change leaves program semantics unchanged, but changing a single keyword can have profound changes. [^1] Compare the instruction “return true if string is float” with “return true if string is verb”. As an example, Elazar et al. (2024) report that only $1.22\%$ of verbatim HumanEval is present in the OSCAR popular web corpus. The same shortcomings of decontamination apply to the creation of large-scale synthetic datasets: for example, the model-generated dataset of Starcoder (Li et al., 2023) is decontaminated only by removing exact docstrings or solutions that match HumanEval or MBPP.
The recent exploration of Riddell et al. (2024) quantifies the proportion of this data leakage in existing datasets using plagiarism tools specifically designed for code. Even when static training datasets are cleaned, contamination may persist through incremental leakage in other sources. For example, entities serving models via an API may encounter these benchmark examples when evaluated by third-party users. When a sample of real model usage is annotated for future training data, samples from benchmark evaluation can leak into future training corpora. Furthermore, these samples may include paraphrases and format changes that further complicate heuristic deduplication. In this scenario, a model may easily memorize completions to purportedly novel prompts. As evidence of this phenomenon, we prompted one popular commercial system (kept anonymous) with partial prompts from HumanEval that were designed to keep the instruction under-specified. Table A in Appendix A shows the outcome and evidence that—despite the ambiguity of the prompt—the result matches exactly the gold solution from HumanEval.
3.2 Data leakage through synthetic data
The most capable of code language models rely heavily on the use of synthetic training data (Xu et al., 2023; Wei et al., 2023, 2024; Llama Team, 2024). A typical pipeline consists of curating prompts related to code generation, inferring completions with a previously trained LLM, and synthesizing unit tests for relevant prompts using LLMs. Completions passing the respective unit tests are considered valid code solutions and can be used as future training examples. Alternatively, if a sufficiently powerful model is used, completions might be used as-is without validation.
Synthetic data unlocks scales that are usually not reachable with human-labeled data. Common synthetic code datasets generally have between tens of thousands and millions of examples. For example, Starcoder2-Instruct is a code dataset of around 238k instances (prior to deduplication) that was created by sampling code from GitHub and using it as seeds to generate self-contained code problems, solutions, and associated tests. evol-instruct is another widely popular dataset used by code LLMs such as WizardCoder (Xu et al., 2023) for training. It comprises $110k$ complex query prompts with non-verified completions from closed and open-source models. [^2] Per downloads, the most popular version is a ‘lightly decontaminated’ version on HuggingFace here. The sheer size of these datasets – compared to the domain they are targeting – might explain some memorization. After all, the number of unique, self-contained interview-like prompts with a reasonable size is fairly limited, and it is possible that synthetic datasets cover a majority of this space. When training data covers most examples of a domain, it does not matter whether the model memorizes the training data or whether it can generalize further. Table 3 shows some examples of very similar (but not necessarily equivalent) data between evol-instruct and MBPP.
However, the use of a synthetic data pipeline might hide real leakage. Prior reports  (Yu et al., 2023, page 8)(Wei et al., 2023, page 4) discuss an apparent high similarity between examples in evol-instruct and HumanEval. We also found a lot of semantically equivalent prompts between these two datasets and displayed some in Table A. We extend this analysis by studying the similarity between embedded representations [^3] Embedded using Cohere embed v3 (Team, 2024). of the prompts of HumanEval and MBPP with nearest neighbors from evol-instruct and Starcoder-Instruct. Fig. 2 highlights a widespread similarity between synthetic training datasets and public evaluation data, while the similarity with LBPP is uniformly lower. This is despite LBPP having a very similar format to MBPP and HumanEval (short prompts asking to solve logic problems). The main difference between MBPP/ HumanEval and LBPP is the difficulty level: LBPP’s questions are generally harder (see Section 4). This inference aligns with observations when fine-tuning a ‘Command R Refresh’ model adding evol-instruct. HumanEval score increases by +9%, MBPP increases by +2%, but the LBPP score is unchanged.
Whether the high similarity of synthetic datasets with public evaluation data is due to synthetic data filling the space of problems similar to HumanEval/ MBPP or more direct leakage (eg, through the use of in-context examples), these results point to a larger issue. HumanEval/ MBPP cannot be used as the only proxies to evaluate a model’s code abilities. They mostly provide code performance signal on a very specific type of problems with a very specific level of difficulty. We need more diversity in the code evaluation benchmarks and we believe that LBPP is a step in the right direction.
3.3 Overfitting to test sets
The exaggerated importance of a few benchmarks encourages an incentive structure where model selection prioritizes gain on a narrow suite of metrics. While it may be tempting to use such benchmarks as a deciding factor between similar checkpoints, section 3.2 shows that the correlation between these benchmarks and ‘solving code generation’ is weak. While the meaning and measurement of this unscientific objective are subject to constant revision, selecting for optimal HumanEval performance is akin to $p$-hacking in other fields. This can be justified by assuming these benchmarks are the new dev sets, while the true test is the usage of users over time. However, the usefulness of a dev set entirely relies on its similarity with the actual use case.
Moreover, some risk remains that models overfit to these ‘lucrative’ benchmarks, distorting the perception of downstream performance. Table 2 and Fig. 3 illustrate this problem well. Even though the correlation between MBPP/HumanEval scores and LBPP scores is strong, some models are ranking noticeably higher on MBPP/HumanEval than on LBPP.
The ultimate effect is imbalanced optimization solely towards these metrics, further motivating the practices outlined in Sections 3.1 and 3.2.
Unsolved problems in LBPP
Given a list of unique words each of size $k$
and an $n$ sized word, w, where n is a multiple of $k$,
write a program in Python to determine the number
of unique combinations of words in the list that can
be concatenated to form an anagram of the word.
Write a function in python that takes as input a
recursive function and some parameters. The function
should return the number of times the recursive function
ran itself when starting it with the provided parameters.
Write a Monte Carlo function in Python to compute
the median number of cards you’d need to draw
from a deck such that the sum equals or exceeds
the value $V$.
Table 1: Sampled prompts in LBPP unsolved by leaders on existing benchmarks. Prompts shortened for brevity.
4 LBPP: Less Basic Python Problems
As mentioned above, we have created Less Basic Python Problems (LBPP) for a less biased measure of modern models’ code capabilities. LBPP is a dataset of 161 code completion problems in the style of HumanEval. All prompts include evaluation unit tests, with a median of 4 tests per example.
Dataset Annotation:
Human annotators were asked to create brand new problems that were not solvable by a strong internal model in the loop. All annotators had competitive programming experience. They were instructed to come up with unique problems either from scratch or inspired by programming textbooks whose content was not freely available (e.g., searchable or unlicensed) on the Internet. Annotators could not copy any exercise from a web source or any LLM and only use these sources for inspiration. All sources were cited by annotators and prompts were verified to not match source inspiration. Every prompt is also manually verified as not easily searchable on the web at the time of writing. Each prompt went through additional review to ensure that they were original, hard, and unambiguous. Around a third of the suggested prompts were disqualified for one of these reasons.
All annotators were paid above minimum wage in their respective countries, and all final prompt-completion pairs were manually reviewed by the authors. This adversarial collection resulted in more difficult problems, with most models solving less than $50\%$ of the dataset.
[…]
Mistral Mistral 7B 0.31 0.32 0.11 27$\rightarrow$26
Mixtral $8\times 7$B 0.53 0.23 0.17 22$\rightarrow$23
Mixtral $8\times 22$B 0.73 0.64 0.38 13$\rightarrow$11
Mistral Large 0.92 0.74 0.50 1$\rightarrow$5
Codestral 22B 0.82 0.48 0.40 5$\rightarrow$9
Meta Codellama 7B Instruct 0.39 0.37 0.14 25$\rightarrow$24
Codellama 34B Instruct 0.53 0.53 0.19 23$\rightarrow$22
Llama2 7B Chat 0.17 0.19 0.02 28$\rightarrow$28
Llama2 60B Chat 0.32 0.31 0.10 26$\rightarrow$27
Llama3 8B Instruct 0.62 0.44 0.27 20$\rightarrow$18
Llama3 70B Instruct 0.82 0.67 0.49 6$\rightarrow$6
OpenAI GPT-3.5 Turbo 0.75 0.70 0.37 9$\rightarrow$12
GPT-4 0.82 0.80 0.53 7$\rightarrow$4
GPT-4o 0.90 0.80 0.63 2$\rightarrow$2
Antropic Claude-2 0.65 0.39 0.34 17$\rightarrow$15
Claude-3-Haiku 0.77 0.64 0.34 8$\rightarrow$14
Claude-3-Sonnet 0.74 0.66 0.40 10$\rightarrow$8
Claude-3-Opus 0.84 0.75 0.54 4$\rightarrow$3
Claude-3.5-Sonnet 0.88 0.78 0.64 3$\rightarrow$1
Qwen Qwen1.5 72B Chat 0.63 0.53 0.20 19$\rightarrow$21
Qwen1.5 110B Chat 0.73 0.57 0.30 12$\rightarrow$16
Qwen 2 72B Instruct 0.74 0.67 0.42 11$\rightarrow$7
Cohere Command R 0.43 0.46 0.12 24$\rightarrow$25
Command R (Refresh) 0.71 0.55 0.35 15$\rightarrow$13
Command R+ 0.65 0.61 0.22 18$\rightarrow$20
Command R+ (Refresh) 0.68 0.61 0.29 16$\rightarrow$17
Deepseek Coder 33B Instr. 0.73 0.66 0.39 14$\rightarrow$10
Databricks DBRX Instr. 0.60 0.57 0.25 21$\rightarrow$19
Table 2: Pass@1 results across popular models for zero-shot HumanEval, MBPP and LBPP. All models perform worse on LBPP than either existing benchmark. Rankings between models change between HumanEval and LBPP, contrasting to Riddell et al. (2021). Model rankings similarly also change between MBPP and LBPP.
Initial Results:
Table 2 shows Pass@1 on LBPP for a range of models. Notably, leading models for HumanEval and MBPP perform up to $43\%$ and $27\%$ poorer on LBPP respectively. Additionally, model rankings between either HumanEval and MBPP update for LBPP, potentially identifying overfitting to public benchmarks when presented with a challenging, unseen test set. LBPP is a similarly reliable indicator of code generation performance than prior benchmarks. In Fig. 3, we observe a strong significant correlation between ‘Pass@1’ scores of either HumanEval or MBPP and LBPP. While existing public benchmarks are still a valuable target signal for performance, LBPP is additionally advantageous in that the problems are harder (see Table 2), the dataset is uncontaminated in current training corpora, and prompts bear lower resemblance to existing synthetic corpora (see Fig. 2).
@@ 4 LBPP: Less Basic Python Problems
Figure 3: Pass@1 rate of LBPP against (a) HumanEval and (b) MBPP. LBPP performance correlates with both prior datasets, but is designed to be genuinely unseen by contemporary LLMs.
Challenges in LBPP:
We study common errors and mistakes from multiple models to identify the most challenging features in LBPP. Examples of problems unsolved by all models are shown in Table 1. Considering common errors between Claude 3.5-Sonnet and Command R Refresh, as the best and recently released model respectively, we identify multiple core trends in failure. Of mutual errors: 21% are related to problems on 2D and 3D arrays; 18% are related to graph-oriented algorithms; and 17% are concerning complex programming concepts often used in competition settings. Additional challenging topics include bit arithmetic & manipulation (8%), Pandas data processing (8%), and file IO (8%). The range of shortcomings between all models highlights the variety of domains that future LLMs must master to improve code generation. Overall, the diversity and difficulty of the problems in LBPP challenges even purportedly advanced models with novel and unsolved prompts.